### v1

@@ 3.2 Referential Ordering using LLM
Take the FE module in $R_{1}$ as an example, as depicted in Fig. 1(b), we start with the lower branch which locally emphasizes the features of the potential anchor/target objects via a cross-attention layer by treating $F^{mask}_{i}$ as value/key and encoded text feature of $O_{1:B}$ (denoted as $T_{O_{1:B}}$) as query. On the other hand, the upper branch of Fig. 1(b) explores the spatial relations between all objects by treating the self-attended $F_{1}$ as key/value of another cross-attention, with the concatenation of $T_{O_{1:B}}$ and $T$ being query. Finally, an additional cross-attention layer is applied to the previous output features of both branches to obtain the enhanced proposal feature $F_{2}$, which enriches not only the information of anchor/target objects but also the relations between them.
@@
where $M_{1}$ represents the $K$-dimensional binary mask as defined in Eq. 1. It is worth noting that, $\mathcal{L}_{mask}$ is applied to output features of each referring block with a similar formulation to ensure each output feature contains the information of the current anchor/target objects correspondingly.
3.4 Order-Aware Pre-training with Synthetic Referential Order
Although DOrA is designed to produce a referential order of anchor objects for localizing the target object, only the ground truth point cloud information of the target object is given during training. Thus, the above framework is viewed as a weakly-supervised learning scheme since there is no ground truth referential order available during training. To provide better training supervision, we pre-train DOrA with a simple yet proper synthetic 3D visual grounding task, where the ground-truth labels of anchor/target objects and descriptions with accurate referential orders can be obtained. This pre-training strategy is presented below.
@@ 3.4.1 Augmenting plausible referential order and description
To provide better training supervision and to ensure the reliability of our synthesized data, the constructed description and the corresponding referential order need to be easily and uniquely determined based on anchor/target objects. In our work, we choose to consider spatial relations between objects that are independent of viewpoint (e.g., “nearest” or “farthest”) as the constructing descriptions, suggesting the referential order of anchor/target objects during this data augmentation stage. As highlighted in Algorithm A1 of our supplementary material, given $P$, $L$, and $B$, we construct an augmented referential order $O^{aug}_{1:B}$ by choosing $B$ different class labels $\{l^{aug}_{1},\cdots,l^{aug}_{B}\}$ from $L$ and extracting their class names. The augmented description $D^{aug}$ is then derived as:
@@
Since $D^{aug}$ is constructed following the appearing sequence of object names in $O^{aug}_{1:B}$, it is guaranteed that $O^{aug}_{1:B}$ is a correct referential order w.r.t. $D^{aug}$ and thus can be served as ground truth supervision for pre-training DOrA.
@@
3.4.2 Pre-training objectives
To have DOrA follow $O^{aug}_{1:B}$ to refer $p^{aug}_{i}$ in the $i$-th referring block $R_{i}$, we design a coordinate loss $\mathcal{L}_{crd}$ to encourage the output feature $F_{i+1}$ of $R_{i}$ to identify the coordinate of all proposals in $P$ w.r.t. $p^{aug}_{i}$. Thus, we calculate $\mathcal{L}_{crd}$ as:
@@ 3.4.2 Pre-training objectives
We note that, the referential loss $\mathcal{L}_{ref}$ mentioned in Sec. 3.3 is also extended to classify both the identity of anchor objects using $F_{2:B}$ and the identity of the target object for $F_{B+1}$ during the pre-training. To this end, we can define the objectives used during our pre-training by summing up the referential loss $\mathcal{L}_{ref}$ (for both anchor and target objects), the masking loss $\mathcal{L}_{mask}$, the language classification loss $\mathcal{L}_{text}$ and the coordinate loss $\mathcal{L}_{crd}$.
@@ 3.5 Overall Training Pipeline
We now summarize the training of DOrA. With the pre-training stage noted in Sec. 3.4 to warm up the model, we take point cloud data with real-world natural descriptions to continue the training process. Since the identity of anchor objects is unknown, we only apply $\mathcal{L}_{ref}$ (for the target object only), $\mathcal{L}_{mask}$, and $\mathcal{L}_{text}$ as supervision. The overall training pipeline is summarized in Algorithm A2 of our supplementary material.
@@ 4 Experiments
4.1 Datasets
4.1.1 NR3D
NR3D [2] dataset consists of 707 indoor scenes in ScanNet [12] with 28715/7485 description-target pairs in the training/testing set, where the descriptions are collected from human annotators. There are 524 different object classes in the scenes in total. NR3D assumes perfect class-agnostic object proposals, where each point in the scene is properly assigned to its corresponding proposal. As a result, models are only required to classify the target object that uniquely matches the description among all proposals in the scene, with classification accuracy (Acc in %) being the metric. Please refer to the supplementary material for the details of the implementation.
4.1.2 ScanRefer
ScanRefer [8] contains 36665/9508 description-target pairs across a total of 800 indoor scenes in its training/validation set, where the descriptions are also collected from human annotators. Also derived from ScanNet [12] but different from NR3D, perfect object proposals are not available in ScanRefer, and therefore additional object proposers are required for all methods. Nevertheless, Acc in % under 0.25 and 0.5 intersection over union (IoU) are used as the metrics for ScanRefer. Please refer to the supplementary material for the details of the implementation.
4.2 Quantitative Results
Table 1: Grounding accuracy (%) on NR3D. Note that each column shows the results trained with a specific amount of training data.
@@ 4.2 Quantitative Results
1% 2.5% 5% 10% 100%
Referit3D [2] 4.4 13.6 20.3 23.3 35.6
TransRefer3D [17] 11.0 16.1 21.9 25.7 42.1
SAT [46] 11.6 16.0 21.4 25.0 49.2
BUTD-DETR [21] 24.2 28.6 31.2 33.3 54.6
MVT [20] 9.9 16.1 21.6 26.5 55.1
MVT [20] + CoT3DRef [5] - - - 38.2 60.4
DOrA (Ours) 33.5 36.1 41.5 46.0 59.7
We present the quantitative results on NR3D and ScanRefer, with consideration of different amounts of available training samples against several baselines by reproducing from their official implementation. As shown in Tab. 1, DOrA possesses a considerably superior performance when training with 1% (287), 2.5% (717), 5% (1435), and 10% (2871) NR3D training set samples. Specifically, using only 10% data, DOrA achieves 46.0 overall Acc, which outperforms Referit3D and TransRefer3D with 100% (28715) data. When training with 100% data, DOrA achieves comparable results to MVT augmented with CoT3DRef [^1] Performance of MVT+CoT3DRef on 1%, 2.5%, and 5% training data are not available since its referential order and 1-to-1 class-identity relation are not publicly available. Only their results of 10% and 100% data are presented., which requires additional prior linguistic knowledge about the dataset for its rule-based algorithm to construct a one-to-one relation between each anchor’s class name and corresponding object identity.
Tab. 2 further displays the detailed performance on different official subsets of NR3D. Among the subsets, the Hard subset contains samples with more than 2 distractors, where a distractor is an object having the same class name as the target object, and the Easy subset is the contrary. View-dependent samples contain relations where rotating the point cloud scene will affect the referred ground-truth target object (e.g., left and right), and View-independent samples are contrary. DOrA accentuates itself with decent capabilities on different subsets, consistently followed by BUTD-DETR [21] that also features a self-designed DETR-based data augmentation for 3D visual grounding.
Finally, Tab. 3 presents DOrA’s performance on ScanRefer, where ground-truth object proposals are unavailable. We can see that our DOrA demonstrates slightly preferable results with only 5% and 10% of data compared to other approaches. It is comparable to SOTA methods when given 100% data. Note that we only report CORE-3DVG’s [44] results on 100% (36665) ScanRefer training samples since its implementation is currently unavailable.
@@
DOrA (Ours) 39.1 53.3 45.3 46.4 46.0
Table 3: Grounding accuracy (%) on the ScanRefer validation set. Note that different amounts of training data are considered.
Method Labeled Training Data
5% 10% 100%
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [8] 23.0 12.0 27.4 15.0 35.5 22.4
3DVG-Trans. [53] 35.3 23.3 39.0 29.0 45.9 34.5
3D-SPS [31] 28.4 16.9 32.9 22.9 48.8 37.0
BUTD-DETR [21] 38.2 26.2 40.3 28.3 52.2 39.8
M3DRef-CLIP [52] 37.2 29.9 40.0 32.4 52.5 44.7
CORE-3DVG [44] - - - - 56.8 43.8
DOrA (Ours) 39.8 31.5 43.6 34.9 52.8 44.8
@@ 4.3 Ablation Studies
4.3.1 Component analysis
We ablate different components of our DOrA in Tab. 4, with performances under 1% and 100% training samples available. Baseline model A extracts class names of anchor/target objects from $D$ using a language parser [6] and constructs the referential order according to the appearance of the names in $D$ directly. Also, we employ each of our Referring Blocks $R_{i}$ in model A without the FE module, i.e., directly using $F_{i}$ to calculate attention with text features of $O_{i:B}$ and $D$. Model B enhances the accuracy on 1% training data by conducting our two-stage object ordering with LLM. When further applying the order-aware pre-training in model C, significant improvements in both settings, especially for 1% available data, are observed. The scarcity of training data often impedes the ability of the model to learn the concept of order and relations among objects implicitly. By conducting our pre-training strategy, our DOrA is able to learn foundational concepts of ordering and relations to locate the target objects progressively. Finally, our full model in the last row, incorporating FE modules each $R_{i}$ to enhance features of anchor/target objects with $F^{mask}_{i}$, achieves optimal results on both settings. This verifies the success of our proposed modules and pre-training strategy.
Table 4: Ablation studies of proposed components in DOrA. For methods without LLM Object Ordering, we form $O_{1:B}$ according to the appearance of anchor/target object names in $D$. Both cases of 1% and 100% training data are considered.
FE Module Order-Aware Pre-training LLM Object Ordering 1% 100%
A 8.9 53.9
B 10.3 53.8
C 25.8 58.0
Ours 33.5 59.7
(a) Grounding accuracies with different maximum (padded) referential order length $B$. Note that $B$ is equal to and denoted as the number of Object-Referring blocks.
Figure 3: (a) Grounding accuracies with varying Object-Referring block numbers $B$ and (b) statistics of referential order length of NR3D. It can be seen that the model performance saturates at $B$=4, matching length statistics of NR3D.
4.3.2 Length of referential order
We additionally analyze the effectiveness of padded order length $B$ of $O_{1:B}$ (and also the number of Referring Blocks) in 3, which illustrates the performances with different $B$ using 1% and 100% training data in Fig. 2(a). We also show the statistics of the original order length generated by our two-stage referential ordering in Fig. 2(b). We observe that although performances gradually increase with larger $B$ (i.e., considering more potential anchor objects appeared in $D$ achieves better accuracy), the performance gain saturates at $B=4$. This can be explained by looking at Fig. 2(b), which shows that only a very few amount of data exceeds an order length of $4$ in both training and testing pairs. This suggests that our selection of $B=4$ is reasonable, optimally balancing computational efficiency with prediction accuracy.
@@ 4.4 Qualitative Results
Fig. 4 demonstrates the qualitative results of DOrA on NR3D, with MVT being the baseline. We display four successful cases and two failed cases. It is shown that DOrA can successfully identify the target object referred by one to multiple anchor objects, even in lengthy descriptions. Failed cases include those that refer by shapes or have a very small target object that is hard for Pointnet++ to capture visual information.
@@
Figure 4: 3D grounding examples of NR3D. Note that blue/green/red boxes denote ground truth/correct/incorrect predictions. While both MVT and DOrA fail on the last two cases, it is due to the fact that the size of the target object is extremely small (e.g., cup) and the description does not describe any anchor objects.
5 Conclusions and Limitations
5.0.1 Conclusions
@@
We presented a 3D visual grounding framework with Order-Aware referring (DOrA) in this paper. DOrA leverages LLM to identify anchor/target objects from the input description as a referential order, guiding the updates of the associated object features for grounding purposes. The above process is realized by stacked Object-Referring blocks in DOrA, which progressively refine the features for the objects of interest in the referential order. In addition, we also designed a pre-training scheme which augments a pseudo yet correct series of anchor/target objects for warming up DOrA in a supervised fashion. Experiments on benchmark datasets demonstrate that our DOrA performed favorably against SOTA 3D grounding works across different settings.
5.0.2 Limitations
Since DOrA utilizes LLMs to generate the referential order from the input description, despite of our introduced pre-training strategy to warm the training process, the correctness of the extracted referential order would affect the training and the performance of DOrA.
Appendix 0.A More Details of DOrA
@@
This section provides more details regarding model implementations and experimental setups. In particular, we first demonstrate the pseudocode of order-aware sample synthesis for pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5. Then, we elaborate on the hyperparameters of DOrA and implementations of baselines for experiments in Sec. 4.
0.A.1 Pre-Training Sample Synthesis and Training Pipeline
@@
This section provides the pseudocode of synthesizing order-aware samples for DOrA’s pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5 in our main paper, respectively. Specifically, the Algorithm A1 demonstrates the pipeline to synthesize an order-aware pre-training sample given simply object proposals $P$ and predicted object labels $L$. On the other hand, Algorithm A2 illustrates the complete pipeline to train DOrA with synthesized samples and natural-description samples, such as NR3D and ScanRefer.
Algorithm A1 Order-Aware Sample synthesis for DOrA Pre-training
@@
0.A.2 Implementation and Details and Experimental Settings
For both datasets, we use PyTorch [33] library to implement DOrA. We train DOrA using Adam [24] optimizer with a single NVIDIA Tesla V100 GPU. To conduct batch-wise training with a fixed number of Object-Referring blocks, we set the length of referential order $B$ to be 4, i.e., we trim the original referential order from the front if its length exceeds 4 and pad it if its length is lower than 4. We adopt BERT [13] as the text encoder to extract $T$ and Pointnet++ [35] as the visual encoder to acquire $F_{1}$. We sample $I$=1024 points for each object proposal in the scene. Object proposal number $K$ and token number $|D|$ are sample-dependent. Object feature dimension $d_{i}$ for $i$-th Object-Referring block is set to 768, aligning with BERT’s 768 dimension, to conduct cross-attention.
Algorithm A2 DOrA Training Pipeline

### v2

@@ 3.2.2 Object feature enhancement
Take the FE module in $R_{1}$ as an example, as depicted in Fig. 3(a), we start with the lower branch which locally emphasizes the features of the potential anchor/target objects via a cross-attention layer by treating $F^{mask}_{i}$ as value/key and encoded text feature of $O_{1:B}$ (denoted as $T_{O_{1:B}}$) as query. On the other hand, the upper branch of Fig. 3(a) explores the spatial relations between all objects by treating the self-attended $F_{1}$ as the key/value of another cross-attention, with the concatenation of $T_{O_{1:B}}$ and $T$ being query. Finally, an additional cross-attention layer is applied to the previous output features of both branches to obtain the enhanced proposal feature $F_{2}$, which enriches not only the information of anchor/target objects but also the relations between them.
@@
where $M_{1}$ represents the $K$-dimensional binary mask as defined in Eqn. 1. It is worth noting that, $\mathcal{L}_{mask}$ is applied to output features of each referring block with a similar formulation to ensure each output feature contains the information of the current anchor/target objects correspondingly.
3.3 Order-Aware Warm-up with Synthetic Referential Order
Although Vigor is designed to produce a referential order of anchor objects for localizing the target object, only the ground truth point cloud information of the target object is given during training. Thus, the above framework is viewed as a weakly-supervised learning scheme since there is no ground truth referential order available during training. To provide better training supervision, we warm-up Vigor with a simple yet proper synthetic 3D visual grounding task, where the ground-truth labels of anchor/target objects and descriptions with accurate referential orders can be obtained. This warm-up strategy is presented below.
@@ 3.3.1 Augmenting plausible referential order and description
To provide better training supervision and to ensure the reliability of our synthesized data, the constructed description and the corresponding referential order need to be easily and uniquely determined based on anchor/target objects. In our work, we choose to consider spatial relations between objects that are independent of viewpoint (e.g., “nearest” or “farthest”) as the constructing descriptions, suggesting the referential order of anchor/target objects during this data augmentation stage. As highlighted in Fig. 3(b) and Algorithm A1 of our supplementary material, given $P$, $L$, and $B$, we construct an augmented referential order $O^{aug}_{1:B}$ by choosing $B$ different class labels $\{l^{aug}_{1},\cdots,l^{aug}_{B}\}$ from $L$ and extracting their class names. The augmented description $D^{aug}$ is then derived as:
@@
Since $D^{aug}$ is constructed following the appearing sequence of object names in $O^{aug}_{1:B}$, it is guaranteed that $O^{aug}_{1:B}$ is a correct referential order w.r.t. $D^{aug}$ and thus can be served as ground truth supervision for pre-training Vigor.
@@
3.3.2 Warm-up objectives
To have Vigor follow $O^{aug}_{1:B}$ to refer $p^{aug}_{i}$ in the $i$-th referring block $R_{i}$, we design a coordinate loss $\mathcal{L}_{crd}$ to encourage the output feature $F_{i+1}$ of $R_{i}$ to identify the coordinate of all proposals in $P$ w.r.t. $p^{aug}_{i}$. Thus, we calculate $\mathcal{L}_{crd}$ as:
@@ 3.3.2 Warm-up objectives
We note that, the referential loss $\mathcal{L}_{ref}$ mentioned in Sec. 3.2 is also extended to classify both the identity of anchor objects using $F_{2:B}$ and the identity of the target object for $F_{B+1}$ during the warm-up process. To this end, we can define the objectives used during our warm-up process by summing up the referential loss $\mathcal{L}_{ref}$ (for both anchor and target objects), the masking loss $\mathcal{L}_{mask}$, the language classification loss $\mathcal{L}_{text}$ and the coordinate loss $\mathcal{L}_{crd}$. With this warm-up strategy, Vigor is initialized to observe relations between anchor/target objects before the subsequent fully-supervised training stage. Later we will verify that, with this pre-training scheme, our Vigor produces satisfactory grounding performances especially when the amount of supervised training data is limited.
@@ 3.4 Overall Training Pipeline
We now summarize the training of Vigor. With the warm-up stage noted in Sec. 3.3, we take point cloud data with real-world natural descriptions to continue the training process. Since the identity of anchor objects is unknown, we only apply $\mathcal{L}_{ref}$ (for the target object only), $\mathcal{L}_{mask}$, and $\mathcal{L}_{text}$ as supervision. The overall training pipeline is summarized in Algorithm A2 in supplementary.
@@ 4 Experiments
4.1 Dataset
NR3D
NR3D [2] dataset consists of 707 indoor scenes in ScanNet [12] with 28715/7485 description-target pairs in the training/testing set, where the descriptions are collected from human annotators. There are 524 different object classes in the scenes in total. NR3D assumes perfect class-agnostic object proposals, where each point in the scene is properly assigned to its corresponding proposal. As a result, models are only required to classify the target object that uniquely matches the description among all proposals in the scene, with classification accuracy (Acc in %) being the metric. Please refer to the supplementary material for the details of the implementation (A) and experiments on an additional dataset ScanRefer [8] B.2.1.
4.2 Quantitative Results for Data Efficiency
Table 1: Data Efficient Grounding accuracy (%) on NR3D. Note that each column shows the results trained with a specific amount of training data.
@@ 4.2 Quantitative Results for Data Efficiency
1% 2.5% 5% 10%
Referit3D [2] 4.4 13.6 20.3 23.3
TransRefer3D [17] 11.0 16.1 21.9 25.7
SAT [46] 11.6 16.0 21.4 25.0
BUTD-DETR [21] 24.2 28.6 31.2 33.3
MVT [20] 9.9 16.1 21.6 26.5
MVT + CoT3DRef [5] 9.4 17.3 26.5 38.2
ViL3DRel + CoT3DRef [5] 22.38 27.25 33.77 38.4
Vigor (Ours) 33.5 36.1 41.5 46.0
We present the quantitative results on NR3D, with consideration of different amounts of available training samples against several baselines by reproducing from their official implementation. As shown in Table 1, Vigor possesses a considerably superior performance when training with 1% (287), 2.5% (717), 5% (1435), and 10% (2871) NR3D training set samples. Specifically, using only 1% data, Vigor achieves 33.5 overall Acc, which even surpasses SOTA methods with 10% of data. This suggests that Vigor is preferable for 3D grounding, especially when paired training data is limited.
Table 2 further displays the detailed performance on different official subsets of NR3D. Among the subsets, the Hard subset contains samples with more than 2 distractors, where a distractor is an object having the same class name as the target object, and the Easy subset is the contrary. View-dependent samples contain relations where rotating the point cloud scene will affect the referred ground-truth target object (e.g., left and right), and View-independent samples are contrary. Vigor accentuates itself with decent capabilities on different subsets, consistently followed by the two variants of CoT3DRef [5] that also feature the concept of referential order for 3D visual grounding.
@@
MVT + CoT3DRef [5] 32.7 43.2 34.0 39.8 37.9
ViL3DRel + CoT3DRef [5] 32.4 44.7 33.4 40.9 38.4
Vigor (Ours) 39.1 53.3 45.3 46.4 46.0
@@ 4.3 Ablation Studies
We ablate different components of our Vigor in Table 3, with performances under 1%, 10%, and 100% training samples available to explore the effectiveness of each component with different amounts of data. Baseline model A extracts class names of anchor/target objects from $D$ using a language parser [6] and constructs the referential order according to the appearance of the names in $D$ directly. Also, we employ each of our Referring Blocks $R_{i}$ in model A without the FE module, i.e., directly using $F_{i}$ to calculate attention with text features of $O_{i:B}$ and $D$. Model B enhances the accuracy on 1% and 10% training data by conducting our two-stage object ordering with LLM. When further applying the order-aware pre-training in model C, significant improvements in all settings, especially for 1% available data, are observed. The scarcity of training data often impedes the ability of the model to learn the concept of order and relations among objects implicitly. By conducting our pre-training strategy, our Vigor is able to learn foundational concepts of ordering and relations to locate the target objects progressively. Finally, our full model in the last row, incorporating FE modules each $R_{i}$ to enhance features of anchor/target objects with $F^{mask}_{i}$, achieves optimal results on both settings. This verifies the success of our proposed modules and warm-up strategy, especially when the available training data is very limited.
Table 3: Ablation studies of proposed components in Vigor. For methods without LLM Object Ordering, we form $O_{1:B}$ according to the appearance of anchor/target object names in $D$. Cases of 1%, 10%, and 100% training data are considered.
FE Module Order-Aware Pre-training LLM Object Ordering 1% 10% 100%
A 8.9 35.2 53.9
B 10.3 38.8 53.8
C 25.8 42.5 58.0
Ours 33.5 46.0 59.7
@@ 4.4 Qualitative Results
Fig. 4 demonstrates the qualitative results of Vigor on NR3D, with MVT being the baseline. We display four successful cases and two failed cases. It is shown that Vigor can successfully identify the target object referred by one to multiple anchor objects, even in lengthy descriptions. Failed cases include those that refer by shapes or have a very small target object that is hard for Pointnet++ to capture visual information.
@@
Figure 4: 3D grounding examples of NR3D. Note that blue/green/red boxes denote ground truth/correct/incorrect predictions. While both MVT and Vigor fail on the last two cases, it is due to the fact that the size of the target object is extremely small (e.g., cup) and the description does not describe any anchor objects.
5 Conclusions
@@
We presented a data-efficient 3D Visual Grounding Framework with Order-Aware Referring (Vigor) in this paper. Vigor identifies anchor/target objects from the LLM-parsed referential order of input description and guides the updates of the associated object features for grounding purposes. The above process is realized by stacked Object-Referring blocks in Vigor, which progressively process the features of the objects of interest in the above referential order. In addition, a unique warm-up scheme to pre-train Vigor was presented, that augments a pseudo yet desirable series of anchor/target objects and enables Vigor to realize relations between objects before formal training. Experiments on benchmark datasets demonstrate that our Vigor performed favorably against SOTA 3D grounding works in a data-efficient manner.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A More Details of Vigor
@@
This section provides more details regarding model implementations and experimental setups. In particular, we first demonstrate the pseudocode of order-aware sample synthesis for pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5. Then, we elaborate on the hyperparameters of Vigor and implementations of baselines for experiments in Sec. 4.
A.1 Pre-Training Sample Synthesis and Training Pipeline
@@
This section provides the pseudocode of synthesizing order-aware samples for Vigor’s pre-training in Sec. 3.4 and the complete training pipeline in Sec. 3.5 in our main paper, respectively. Specifically, Algorithm A1 demonstrates the pipeline to synthesize an order-aware pre-training sample given object proposals $P$ and predicted object labels $L$. On the other hand, A2 illustrates the complete pipeline to train Vigor with synthesized samples and natural-description samples, such as NR3D and ScanRefer.
Algorithm A1 Order-Aware Sample synthesis for Vigor Pre-training
@@
A.2 Implementation and Details and Experimental Settings
For both datasets, we use PyTorch [33] library to implement Vigor. We train Vigor using Adam [24] optimizer with a single NVIDIA Tesla V100 GPU.
To conduct batch-wise training with a fixed number of Object-Referring blocks, we set the length of referential order $B$ to be 4, i.e., we trim the original referential order from the front if its length exceeds 4 and pad it if its length is lower than 4. We adopt BERT [13] as the text encoder to extract $T$ and Pointnet++ [35] as the visual encoder to acquire $F_{1}$. We sample $I$=1024 points for each object proposal in the scene. Object proposal number $K$ and token number $|D|$ are sample-dependent. Object feature dimension $d_{i}$ for $i$-th Object-Referring block is set to 768, aligning with BERT’s 768 dimension, to conduct cross-attention.
Algorithm A2 Vigor Training Pipeline